A planet's moon travels in an approximately circular orbit of radius 6.6 107 m with a period of 7 h 19 min. Calculate the mass of the planet from this information. kg
Answer #1

the time period of a satellite around a planet is given by


T = 2\pi \sqrt{\frac{r^{3}}{G*M}}

M is the mass of planet

T = 7 * 3600 +19 * 60 = 26340 s

r = 6.6 * 10^7 m

M = ?

M = 2.45 * 1026 kg